Questions 1 and 2 are based on the following information:

The demand and total cost functions for a commodity are

 

Q = 6 000 − 30P      and     TC = 5 000 + 2Q,

where P and Q are the price and quantity, respectively.

 1

The total revenue function (TR), in terms of P and the company's quantity, if a commodity sells for R199.00                     are?

a. TR = 6 000 − 30P ; Q = 30.

b. TR = 6 000P − 30P 2; Q = 30.

c. TR = P (6 000 − 30P ); Q = 24.

d.  TR = 6 000 − 30P ; Q = 24.

 

2

What is the company's profit function in terms P?

a.   Profit = −30P 2 + 6 060P − 17 000

b.    Profit = −30P 2 + 5 940P − 17 000

c.   Profit = −30P 2 + 5 940P + 7 000

d.    Profit = −30P 2 + 6 060P − 7 000

 

 


1
Expert's answer
2021-06-12T03:40:45-0400

1 The total revenue function (TR), in terms of P and the company's quantity, if a commodity sells for R199.00 are?

TR = Q * P = 6 000 P - 30 P2; Q = 6 000 - 30 P = 6 000 - 30 *199 = 30

Aswer: b


2 What is the company's profit function in terms P?

TC = 5 000 + 2 Q = 5 000 + 2(6 000 - 30 P) = 17 000 - 60 P

Profit = TR - TC = 6 000 P - 30 P2 - (17 000 - 60 P) = -30 P2 + 6 060 P - 17 000

Aswer: a